XYZ's stock price and dividend history are as follows:

Year Beginning-of-Year Price Dividend Paid at Year-End
2017 $ 150 $ 4
2018 180 4
2019 130 4
2020 150 4

An investor buys three shares of XYZ at the beginning of 2017, buys another two shares at the beginning of 2018, sells one share at the beginning of 2019, and sells all four remaining shares at the beginning of 2020.

a. What are the arithmetic and geometric average time-weighted rates of return for the investor? (Round your year-by-year rates of return and final answer to 2 decimal places. Do not round other calculations.)

b. What is the dollar-weighted rate of return? (Hint: Carefully prepare a chart of cash flows for the four dates corresponding to the turns of the year for January 1, 2017, to January 1, 2020. If your calculator cannot calculate IRR, you will have to use trial and error or a spreadsheet program.) (Round your answers to 4 decimal places. Negative amount should be indicated by a minus sign.)
Answer #1
S.No Year Share value at year beginning Dividend No of shares bought No of shares sold Return computation Return %
1 2017                                  150.00 4 3 (Year beginning) =((180-150)+4)/150*100 23%
2 2018                                  180.00 4 2 (Year beginning) =((130-180)+4)/180*100 -26%
3 2019                                  130.00 4 1 (Year beginning) =((150-130)+4)/130*100 18%
4 2020                                  150.00 4 4 (Year beginning)
Return computation formula= ((Year ending price+dividend for the year-year beginning price)/year beginning price)*100
Arithmetic mean of the returns= (23-26+18)/3= 5%
For Geometric mean of the returns please refer to the Photo
Dollar-weighted rate of return
Cash flow
S.No Year Share value at year begining Dividend No of shares bought No of shares sold Year completed Outflow computation Outflow Inflow computation Inflow Net cash flows
1 2017                                  150.00 4 3 (Year beginning)                                   -   =150*3 450 -450
2 2018                                  180.00 4 2 (Year beginning)                              1.00 =180*2 360 =4*3 12 -348
3 2019                                  130.00 4 1 (Year beginning)                              2.00 =(4*5)+(130*1) 150 150
4 2020                                  150.00 4 4 (Year beginning)                              3.00 =(4*4)+(150*4) 616 616
IRR for the investment made in year 1
Year Year 0 investment No of shares Inflow computation Inflow Discounting factor @ Discounted values
0                            450.00 3
1 =4*3                                       12                         0.9901                            11.88
2 =(4*3)+130                                     142                         0.9803                         139.20
3 =4*2+(150*2)                                     308                         0.9705                         298.92
                        450.00
IRR %                         1.0021
IRR for the investment made in year 2
Year Investment made at the end of year 1 No of shares Inflow computation Inflow Discounting factor @ IRR Discounted values
1                            360.00 2 0 0
2 =4*2 8 1.068215941                              8.55
3 =(4*2)+(150*2) 308 1.141085297                         351.45
                        360.00
IRR %                        -6.3860

Geometric mean - | Ct 023) (1-0.26) 40 18) - I024093- O.024013) 2.41/ geametric mean of retums 17